OpenStudy (anonymous):

@whatdoesthismean A student needs to make a square cardboard piece. The cardboard should have a perimeter equal to at least 92 inches. The function f(s) relates the perimeter of a cardboard piece, in inches, to the length of its side in inches. Which of the following shows a reasonable domain for f(s)? 23 < s < 46 23 < s < 92 s ≥ 92 s ≥ 23

OpenStudy (anonymous):

Ok so it is AT LEAST 92. At least always means\[\ge \]

OpenStudy (anonymous):

SO the only one would be C

OpenStudy (anonymous):

Thats what I got also

OpenStudy (anonymous):

Maggie has $32 in her money box. She plans to put $9 in her money box every month. If f(n) represents the total amount of money Maggie will have in her money box after n months, which of the following functions represents the relationship between f(n) and n? (1 point) f(n) = 9n + 32 f(n) = 32n + 9 f(n) = 41n + 9 f(n) = 9n + 41

OpenStudy (anonymous):

But then don't forget that for the last question its regarding only one side, so its D Sorry i did not catch that

OpenStudy (anonymous):

Ok so its not \[s \ge 92\]

OpenStudy (anonymous):

n will always represent time 9 will be the coefficient because she is adding 9 per month, and 32 will be the constant because she already has that and NO it is not \[s \ge92\] it is D

OpenStudy (anonymous):

Alrighty

OpenStudy (anonymous):

Maggie has $32 in her money box. She plans to put $9 in her money box every month. If f(n) represents the total amount of money Maggie will have in her money box after n months, which of the following functions represents the relationship between f(n) and n? f(n) = 9n + 32 f(n) = 32n + 9 f(n) = 41n + 9 f(n) = 9n + 41

OpenStudy (anonymous):

So for the other question, to what I said above, what equation would be built

OpenStudy (anonymous):

n will always represent time 9 will be the coefficient because she is adding 9 per month, and 32 will be the constant because she already has that

OpenStudy (anonymous):

That makes sense

OpenStudy (anonymous):



OpenStudy (anonymous):

So A i think

OpenStudy (anonymous):

Ok awesome